OpenStudy (-welp-):

Algebra 2 - Simplifying Radical Expressions

OpenStudy (-welp-):

\[30\sqrt{50} * \sqrt{22}\] = \[30\sqrt{11}\]

OpenStudy (-welp-):

How did they get this?